# e5-omni: Explicit Cross-modal Alignment for Omni-modal Embeddings

### v1

@@ 2.3 Controllable Negative Curriculum
Tevatron-Omni Ma et al. (2025) 7B 39.2 34.3 68.8 48.1
LCO-EMB Xiao et al. (2025) 7B 45.6 37.0 64.6 50.6
Omni-Embed-Nemotron Xu et al. (2025b) 3B 44.4 35.8 68.6 51.5
@@
e5-omni-vanilla-3B 3B 67.4 39.3 66.7 62.0
e5-omni-vanilla-7B 7B 70.8 40.9 69.0 64.4
@@ Appendix A Detailed Results on MMEB-V2
GME VLM2Vec-v1 VLM2Vec-v2 CAFe UME-R1 LCO-EMB Omni-Embed-Nemotron e5-omni e5-omni Model Size 7B 7B 2B 7B 7B 7B 3B 3B 7B Avg - All (78 tasks) 57.8 52.3 58.0 60.6 64.5 50.6 51.5 63.1 66.4 Avg - Image (36 tasks, Hit@1) 56.0 65.5 64.9 67.6 71.3 45.6 44.4 69.2 73.0 Avg - Video (18 tasks, Hit@1) 38.4 33.7 34.6 42.4 47.5 37.0 35.8 40.6 42.6 Avg - Visdoc (24 tasks, NDCG@5) 75.2 46.4 65.4 63.9 67.1 64.6 68.6 67.9 70.4 I-CLS (10) 57.7 62.7 62.9 63.6 67.1 56.3 46.0 64.7 66.7 I-QA (10) 34.7 56.9 56.3 61.7 69.2 16.9 20.5 62.1 68.5 I-RET (12) 71.2 69.4 69.5 69.1 71.9 52.2 58.0 70.5 73.0 I-VG (4) 59.3 82.2 77.3 87.6 84.9 57.0 52.9 79.6 83.9 V-CLS (5) 37.4 39.1 39.3 35.8 48.6 39.3 40.5 40.2 46.6 V-QA (5) 50.4 30.0 34.3 58.7 60.7 57.6 44.3 48.5 52.9 V-RET (5) 28.4 29.0 28.8 34.4 38.2 24.8 32.7 33.1 36.7 V-MR (3) 37.0 38.9 36.8 39.5 39.3 26.5 25.6 40.7 34.2 VD-Vidore-V1 (10) 89.4 56.9 75.5 70.7 75.7 80.4 85.7 83.1 87.6 VD-Vidore-V2 (4) 55.6 9.4 44.9 49.6 50.5 56.4 61.3 59.5 62.4 VD-VisRAG (6) 85.0 59.1 79.4 79.5 83.7 79.7 84.3 85.5 87.5 VD-OOD (4) 44.4 38.1 39.4 38.1 37.6 41.8 43.3 43.4 44.3 ImageNet-1K 64.6 80.1 80.8 77.3 80.4 73.0 57.0 76.5 79.4 N24News 50.5 79.7 72.9 83.2 82.3 37.5 36.4 76.7 77.2 HatefulMemes 53.6 69.7 56.3 78.7 79.0 45.1 48.0 66.2 72.3 VOC2007 80.3 80.7 85.0 89.8 90.8 64.9 47.9 84.8 83.9 SUN397 69.5 77.4 71.0 79.9 80.3 72.0 53.1 75.1 76.1 Place365 39.1 37.4 35.9 45.0 46.8 40.8 30.7 41.7 44.3 ImageNet-A 41.2 58.1 47.4 55.2 53.9 52.2 37.0 58.6 57.2 ImageNet-R 83.9 73.9 89.3 88.0 90.1 88.9 80.9 85.5 84.5 ObjectNet 69.0 40.1 65.2 22.5 42.3 61.7 44.9 53.0 57.7 Country211 24.8 29.8 25.2 16.7 25.0 26.5 24.1 29.1 34.7 OK-VQA 33.2 56.8 51.5 67.3 71.7 14.4 18.6 64.9 70.5 A-OKVQA 21.0 47.3 43.6 63.8 58.7 10.9 11.9 54.8 59.7 DocVQA 41.4 89.7 90.1 79.2 93.8 17.5 17.3 91.2 93.3 InfographicsVQA 20.3 60.0 58.8 53.3 79.2 7.4 9.1 52.8 62.9 ChartQA 17.8 56.9 47.4 48.8 75.1 12.7 13.1 52.7 62.8 Visual7W 22.2 52.7 52.9 52.5 55.2 10.8 7.4 59.8 65.2 ScienceQA 28.0 38.5 38.2 65.4 53.7 26.7 25.3 43.6 56.4 VizWiz 39.0 39.9 43.3 43.8 51.6 31.8 34.8 53.1 56.3 GQA 76.9 55.1 64.9 65.7 69.3 11.5 31.2 65.4 72.1 TextVQA 46.8 71.6 72.2 76.8 83.5 25.1 36.6 82.8 86.2 VisDial 60.8 81.9 82.7 82.7 80.7 51.1 51.5 78.9 82.4 CIRR 54.9 51.1 57.5 60.4 55.3 8.4 14.8 54.9 54.9 VisualNews_t2i 79.7 80.5 74.5 69.5 76.8 61.4 58.8 74.6 75.3 VisualNews_i2t 83.6 81.2 78.2 79.4 82.0 59.5 56.3 77.7 79.8 MSCOCO_t2i 71.2 77.2 75.3 75.4 78.3 66.1 60.7 78.1 76.6 MSCOCO_i2t 57.7 73.9 71.4 73.1 71.4 59.5 56.5 72.5 76.9 NIGHTS 67.6 67.6 68.6 66.7 68.1 65.2 65.3 68.4 68.6 WebQA 91.4 88.3 90.6 89.3 90.9 53.2 89.8 89.2 89.3 FashionIQ 37.8 17.1 19.5 39.0 23.4 3.5 6.2 19.5 24.5 Wiki-SS-NQ 78.2 62.3 66.9 61.2 72.5 64.2 87.7 74.9 79.4 OVEN 75.1 66.5 64.3 60.8 71.4 59.9 67.9 76.1 78.7 EDIS 96.0 85.7 84.1 71.3 92.0 74.4 80.5 81.7 89.6 MSCOCO 31.4 75.7 67.1 84.7 72.7 33.8 32.8 70.9 74.8 RefCOCO 60.9 87.6 87.1 89.4 91.4 50.9 59.7 85.3 89.4 RefCOCO-Matching 78.4 84.6 85.8 83.0 91.1 98.5 65.9 75.8 87.2 Visual7W-Pointing 66.5 81.0 69.2 93.2 84.2 44.6 53.1 86.5 84.3

### v2

@@ 2.3 Controllable Negative Curriculum
Tevatron-Omni Ma et al. (2025) 7B 37.1 35.1 74.5 48.1
LCO-EMB Xiao et al. (2025) 7B 44.0 38.2 69.8 50.6
Omni-Embed-Nemotron Xu et al. (2025b) 3B 43.7 36.9 74.2 51.5
@@
e5-omni-vanilla-3B 3B 65.8 39.3 72.0 62.0
e5-omni-vanilla-7B 7B 69.0 41.8 74.7 64.4
@@ Appendix A Detailed Results on MMEB-V2
GME VLM2Vec-v1 VLM2Vec-v2 CAFe UME-R1 LCO-EMB Omni-Embed-Nemotron e5-omni e5-omni Model Size 7B 7B 2B 7B 7B 7B 3B 3B 7B Avg - All (78 tasks) 57.8 52.3 58.0 60.6 64.5 50.6 51.5 63.1 66.4 Avg - Image (36 tasks, Hit@1) 56.0 65.5 64.9 67.6 71.3 44.0 43.7 67.6 71.2 Avg - Video (18 tasks, Hit@1) 38.4 33.7 34.6 42.4 47.5 38.2 36.9 40.6 43.5 Avg - Visdoc (24 tasks, NDCG@5) 75.2 46.4 65.4 63.9 67.1 69.8 74.2 73.2 76.1 I-CLS (10) 57.7 62.7 62.9 63.6 67.1 56.3 46.0 64.7 66.7 I-QA (10) 34.7 56.9 56.3 61.7 69.2 16.9 20.5 62.1 68.5 I-RET (12) 71.2 69.4 69.5 69.1 71.9 52.2 58.0 70.5 73.0 I-VG (4) 59.3 82.2 77.3 87.6 84.9 57.0 52.9 79.6 83.9 V-CLS (5) 37.4 39.1 39.3 35.8 48.6 39.3 40.5 40.2 46.6 V-QA (5) 50.4 30.0 34.3 58.7 60.7 57.6 44.3 48.5 52.9 V-RET (5) 28.4 29.0 28.8 34.4 38.2 24.8 32.7 33.1 36.7 V-MR (3) 37.0 38.9 36.8 39.5 39.3 26.5 25.6 40.7 34.2 VD-Vidore-V1 (10) 89.4 56.9 75.5 70.7 75.7 80.4 85.7 83.1 87.6 VD-Vidore-V2 (4) 55.6 9.4 44.9 49.6 50.5 56.4 61.3 59.5 62.4 VD-VisRAG (6) 85.0 59.1 79.4 79.5 83.7 79.7 84.3 85.5 87.5 VD-OOD (4) 44.4 38.1 39.4 38.1 37.6 41.8 43.3 43.4 44.3 ImageNet-1K 64.6 80.1 80.8 77.3 80.4 73.0 57.0 76.5 79.4 N24News 50.5 79.7 72.9 83.2 82.3 37.5 36.4 76.7 77.2 HatefulMemes 53.6 69.7 56.3 78.7 79.0 45.1 48.0 66.2 72.3 VOC2007 80.3 80.7 85.0 89.8 90.8 64.9 47.9 84.8 83.9 SUN397 69.5 77.4 71.0 79.9 80.3 72.0 53.1 75.1 76.1 Place365 39.1 37.4 35.9 45.0 46.8 40.8 30.7 41.7 44.3 ImageNet-A 41.2 58.1 47.4 55.2 53.9 52.2 37.0 58.6 57.2 ImageNet-R 83.9 73.9 89.3 88.0 90.1 88.9 80.9 85.5 84.5 ObjectNet 69.0 40.1 65.2 22.5 42.3 61.7 44.9 53.0 57.7 Country211 24.8 29.8 25.2 16.7 25.0 26.5 24.1 29.1 34.7 OK-VQA 33.2 56.8 51.5 67.3 71.7 14.4 18.6 64.9 70.5 A-OKVQA 21.0 47.3 43.6 63.8 58.7 10.9 11.9 54.8 59.7 DocVQA 41.4 89.7 90.1 79.2 93.8 17.5 17.3 91.2 93.3 InfographicsVQA 20.3 60.0 58.8 53.3 79.2 7.4 9.1 52.8 62.9 ChartQA 17.8 56.9 47.4 48.8 75.1 12.7 13.1 52.7 62.8 Visual7W 22.2 52.7 52.9 52.5 55.2 10.8 7.4 59.8 65.2 ScienceQA 28.0 38.5 38.2 65.4 53.7 26.7 25.3 43.6 56.4 VizWiz 39.0 39.9 43.3 43.8 51.6 31.8 34.8 53.1 56.3 GQA 76.9 55.1 64.9 65.7 69.3 11.5 31.2 65.4 72.1 TextVQA 46.8 71.6 72.2 76.8 83.5 25.1 36.6 82.8 86.2 VisDial 60.8 81.9 82.7 82.7 80.7 51.1 51.5 78.9 82.4 CIRR 54.9 51.1 57.5 60.4 55.3 8.4 14.8 54.9 54.9 VisualNews_t2i 79.7 80.5 74.5 69.5 76.8 61.4 58.8 74.6 75.3 VisualNews_i2t 83.6 81.2 78.2 79.4 82.0 59.5 56.3 77.7 79.8 MSCOCO_t2i 71.2 77.2 75.3 75.4 78.3 66.1 60.7 78.1 76.6 MSCOCO_i2t 57.7 73.9 71.4 73.1 71.4 59.5 56.5 72.5 76.9 NIGHTS 67.6 67.6 68.6 66.7 68.1 65.2 65.3 68.4 68.6 WebQA 91.4 88.3 90.6 89.3 90.9 53.2 89.8 89.2 89.3 FashionIQ 37.8 17.1 19.5 39.0 23.4 3.5 6.2 19.5 24.5 Wiki-SS-NQ 78.2 62.3 66.9 61.2 72.5 64.2 87.7 74.9 79.4 OVEN 75.1 66.5 64.3 60.8 71.4 59.9 67.9 76.1 78.7 EDIS 96.0 85.7 84.1 71.3 92.0 74.4 80.5 81.7 89.6 MSCOCO 31.4 75.7 67.1 84.7 72.7 33.8 32.8 70.9 74.8 RefCOCO 60.9 87.6 87.1 89.4 91.4 50.9 59.7 85.3 89.4 RefCOCO-Matching 78.4 84.6 85.8 83.0 91.1 98.5 65.9 75.8 87.2 Visual7W-Pointing 66.5 81.0 69.2 93.2 84.2 44.6 53.1 86.5 84.3